# Can Dark Stars account for the star formation efficiency excess at very high redshifts?

### v2

@@ 1 Introduction
Dark stars, which inhabit the first dark matter halos or mini halos in the high-redshift universe, are fueled by heating from dark matter (Freese et al., 2008, 2016; Wu et al., 2022; Iocco & Visinelli, 2024; Zhang et al., 2024). This heating may originate from the gravitational attraction of dark matter and the annihilation of dark matter particles, such as weakly interacting massive particles (WIMPs), which can be captured through elastic scattering with baryonic matter (Freese et al., 2010). In the early Universe, dark matter densities were enhanced by a factor of $(1+z)^{3}$ at redshift $z$, suggesting a condition in favor the formation of dark stars (Freese et al., 2016). However, before the launch of JWST, identifying dark stars was challenging due to the lack of very high-redshift observations. Moreover, the outer atmospheric properties of dark stars show similar characteristic compared with those of stars undergoing nuclear reactions, enhancing the difficulty of identification. The dark star observational features might include supermassive dark stars (Iocco et al., 2008; Natarajan et al., 2009; Zackrisson et al., 2010a, b; Ilie et al., 2012; Freese et al., 2016), the extragalactic infrared background light (Maurer et al., 2012), the extragalactic gamma-ray background (Sandick et al., 2011; Yuan et al., 2011; Sandick et al., 2012), remnant black holes following the demise of dark stars (Ilie et al., 2023), the influence of dark stars on the universe’s reionization process (Schleicher et al., 2009; Scott et al., 2011; Gondolo et al., 2022; Qin et al., 2024) and so on.
Recently, three potential supermassive dark star candidates have been identified by JWST (Ilie et al., 2023), for which the power source is rooted in the annihilation of dark matter particles rather than nuclear fusion (Freese et al., 2016). Supposing supermassive dark stars exist indeed, a population of dark stars living within the very high-redshift galaxies could influence the UV LF of galaxies. If the light output of these dark stars can be comparable with normal stars, the formation and evolution of galaxies will be affected in the early universe. In this work, we investigate the possible net contribution of the dark stars to the overall SFE.
@@ 2.2 SFE model
However, dark stars are thought to be powered by dark matter annihilation at the galaxy’s center. Therefore, their formation efficiency likely to be correlated with the mass of the dark matter halo strongly. Compared with normal stars, the formation efficiency of dark stars is more reliant on the density of dark matter, which is directly linked to the mass of the dark matter halo. Given the absence of mature simulations or analytical theories on this topic, it becomes necessary to define a new formation efficiency function specifically for dark stars. For characterizing the formation efficiency of dark stars, we utilize a power-law model expressed as:
@@ 4 Discussion
Note that some interesting models have been proposed to account for the observed excesses in SFE. These include the role of early dark energy (Wang et al., 2024b; Adil et al., 2023) and dynamical dark energy Menci et al. (2024), which might have influenced the expansion rate of the universe and thus the formation of structures. Primordial black holes (Yuan et al., 2024; Su et al., 2023) are another candidate that could contribute to the gravitational potential wells necessary for star formation. Furthermore, the nature of dark matter itself is a subject of ongoing debate (Nesti et al., 2023). Warm dark matter (Lin et al., 2024) and ultra-light dark matter (Gong et al., 2023; Bird et al., 2024) are alternatives to the cold dark matter paradigm, potentially affecting the distribution and dynamics of matter in the early universe. Cosmic strings (Jiao et al., 2023; Wang et al., 2023b), relics from phase transitions in the early universe, could also play a role in structure formation. In addition to these, other mechanisms such as feedback-free processes (Dekel et al., 2023; Li et al., 2024) and Population III star formation (Yung et al., 2024; Inayoshi et al., 2022) should be considered. These mechanisms could provide alternative explanations for the high-redshift SFE excesses and should be distinguished from the new physics mentioned above in future studies.
@@ 5 Acknowledgments
We thank the reviewers for their helpful comments. We thank Steven Finkelstein, Qiao Li, Yong-Jia Huang, Tian-Peng Tang, Yue-Lin Sming Tsai, Hao-Jing Yan, Hai-Bo Yu, Qiang Yuan, Chi Zhang, Tian-Ci Zheng and Hao Zhou for their helpful discussions. This work is supported by the National Key Research and Development Program of China (No.2022YFF0503304), the Natural Science Foundation of China (No.11921003), the China Postdoctoral Science Foundation (No.2023TQ0355), and the New Cornerstone Science Foundation through the XPLORER PRIZE. G.-W. Yuan also acknowledges support from the University of Trento and the Provincia Autonoma di Trento through the UniTrento Internal Call for Research 2023 grant “Searching for Dark Energy off the beaten track” (DARKTRACK, grant agreement no.E63C22000500003, PI: Sunny Vagnozzi).

### v3

@@ 1 Introduction
Dark stars, which inhabit the first dark matter halos or mini halos in the high-redshift universe, are fueled by heating from dark matter (Iocco et al., 2008; Yoon et al., 2008; Freese et al., 2008; Smith et al., 2012; Freese et al., 2016; Wu et al., 2022; Iocco & Visinelli, 2024; Zhang et al., 2024). This heating may originate from the gravitational attraction of dark matter and the annihilation of dark matter particles, particularly weakly interacting massive particles (WIMPs). These particles can potentially be captured through elastic scattering with baryonic matter (Iocco, 2010; Freese et al., 2010) and might be detectable with telescopic observations (Zackrisson et al., 2010a, b) . In the early Universe, dark matter densities were enhanced by a factor of $(1+z)^{3}$ at redshift $z$, suggesting a condition in favor of the formation of dark stars (Freese et al., 2016). However, before the launch of JWST, identifying dark stars was challenging due to the lack of very high-redshift observations. Moreover, the outer atmospheric properties of dark stars show similar characteristics compared with those of stars undergoing nuclear reactions, enhancing the difficulty of identification. The dark star observational features might include supermassive dark stars (Iocco et al., 2008; Natarajan et al., 2009; Zackrisson et al., 2010b, a; Ilie et al., 2012; Freese et al., 2016), the extragalactic infrared background light (Maurer et al., 2012), the extragalactic gamma-ray background (Sandick et al., 2011; Yuan et al., 2011; Sandick et al., 2012), remnant black holes following the demise of dark stars (Ilie et al., 2023), the influence of dark stars on the universe’s reionization process (Schleicher et al., 2009; Scott et al., 2011; Gondolo et al., 2022; Qin et al., 2024) and so on.
Recently, three potential supermassive dark star candidates have been identified by JWST (Ilie et al., 2023), for which the power source is rooted in the annihilation of dark matter particles rather than nuclear fusion (Freese et al., 2016). Supposing supermassive dark stars exist indeed, a population of dark stars living within the very high-redshift galaxies could influence the UV LF of galaxies. If the light output of these dark stars can be comparable with normal stars, the formation and evolution of galaxies will be affected in the early universe. The study by Iocco & Visinelli (2024) estimated the UV limiting magnitudes for a massive dark star, approximately $10^{6}\,\rm M_{\odot}$ in size, situated within a dark matter halo having a mass of $M_{h}=10^{8}\,\rm M_{\odot}$ with a temperature $T=3.14\times 10^{4}\,\rm K$, and an accretion rate of $\dot{M_{h}}=10^{-1}\,\rm M_{\odot}\,yr^{-1}$. Their research concluded that the luminosity of the dar star is insufficient to account for the galaxy’s brightness observed in JWST UV LF studies at $z\sim 13$. In this work, we further investigate the potential contribution of the dark stars to the overall SFE.
@@ 2.2 SFE model
However, dark stars are thought to be powered by dark matter annihilation at the galaxy’s center. Therefore, their formation efficiency is likely to be correlated with the mass of the dark matter halo strongly. Compared with normal stars, the formation efficiency of dark stars is more reliant on the density of dark matter, which is directly linked to the mass of the dark matter halo. Given the absence of mature simulations or analytical theories on this topic, it becomes necessary to define a new formation efficiency function specifically for dark stars. For characterizing the formation efficiency of dark stars, we utilize a power-law model expressed as:
@@ 4 Discussion
Note that some interesting models have been proposed to account for the observed excesses in SFE. These include the role of early dark energy (Wang et al., 2024b; Adil et al., 2023) and dynamical dark energy Menci et al. (2024), which might have influenced the expansion rate of the universe and thus the formation of structures. Primordial black holes (Yuan et al., 2024; Su et al., 2023) are another candidate that could contribute to the gravitational potential wells necessary for star formation. Furthermore, the nature of dark matter itself is a subject of ongoing debate (Nesti et al., 2023). Warm dark matter (Lin et al., 2024) and ultra-light dark matter (Gong et al., 2023; Bird et al., 2024) are alternatives to the cold dark matter paradigm, potentially affecting the distribution and dynamics of matter in the early universe. Cosmic strings (Jiao et al., 2023; Wang et al., 2023b), relics from phase transitions in the early universe, could also play a role in structure formation. In addition to these new physics scenarios, other conventional astrophysical mechanisms should not be overlooked, such as feedback-free processes (Dekel et al., 2023; Li et al., 2024) and Population III star formation (Yung et al., 2024; Inayoshi et al., 2022). Distinguishing between these possibilities will require further theoretical modeling and more precise observational data.
Iocco & Visinelli (2024) demonstrated that the dark stars with masses as large as $10^{6}\,\rm M_{\odot}$ at $z\sim 13$ could in principle generate significant UV luminosity. However, they concluded that such objects are unlikely to explain the brightness JWST UV LFs due to the need for extremely high dark matter density and accretion rate. In this work, we have used the observed spectrum of the $z\sim 13$ galaxy JADES-GS-z13-0, combined with the Hertzsprung-Russell characteristics of dark stars, to constrain the optimal temperature and luminosity contribution of a hypothetical dark star population. Our fitting analysis favors higher dark star surface temperatures, implying either more massive dark stars or heavier WIMPs. Nevertheless, when compared against existing MACHO constraints, we find that the number of black holes required to explain the JWST UV LFs via dark stars would exceed observational limits. This strongly disfavors dark stars as the dominant origin of the UV luminosity observed at the highest redshifts.
@@ 5 Acknowledgments
We thank the reviewers for their helpful comments. We also thank Fabio Iocco, Steven Finkelstein, Qiao Li, Yong-Jia Huang, Tian-Peng Tang, Yue-Lin Sming Tsai, Luca Visinelli, Hao-Jing Yan, Hai-Bo Yu, Qiang Yuan, Chi Zhang, Tian-Ci Zheng and Hao Zhou for their helpful discussions. This work is supported by the National Key Research and Development Program of China (No.2022YFF0503304), the Natural Science Foundation of China (No.11921003), the China Postdoctoral Science Foundation (No.2023TQ0355), and the New Cornerstone Science Foundation through the XPLORER PRIZE. G.-W. Yuan also acknowledges support from the University of Trento and the Provincia Autonoma di Trento through the UniTrento Internal Call for Research 2023 grant “Searching for Dark Energy off the beaten track” (DARKTRACK, grant agreement no.E63C22000500003, PI: Sunny Vagnozzi).
@@ Appendix B SFE result of UV Luminosity Function Fitting
We also plot the corners of the posteriors in Figures 5 and 6. The dark red regions of the three depths in the corner plots of 5 and 6 represent 68%, 95%, and 99% confidence intervals, respectively. The black crosses are the positions of the best-fit parameters.